OpenStudy (anonymous):

how do I figure out sec(theta)=-2?

jimthompson5910 (jim_thompson5910):

sec(theta)=-2 1/cos(theta)=-2 1=-2*cos(theta) 1/(-2) = cos(theta) cos(theta) = -1/2 Does that help?

OpenStudy (anonymous):

yes thank you!

jimthompson5910 (jim_thompson5910):

yw